Beam Currents observed to be consistent today. PM downtime starts at 12 NOON.
jwc at snf.stanford.edu
Thu Feb 12 13:22:02 PST 2004
I was unable to recreate the circumstances that resulted in the erratic
beam currents that were reported by users yesterday.
MEASURED IN SPOT MODE ON THE STAGE FARADAY CUP:
30 um aperture == 0.151 nA 30 measurements
10 um aperture == 0.020 na. 10 measurements
This was measured repeatable over a two hour period and no changes were
observed. All seems good on the system from these test and imaging on
the LEO was excellent resolving features to about 12 nm.
Tomorrow Michael Santomango from LEO will arrive to perform a complete
annual PM on the LEO column about noon.
Users are asked to yield access to him upon his arrival and assist him
any way you can to complete the PM.
We will then bakeout the column for 40 hours and through the entire
Tuesday we will realign the column, mag. amplifiers, stigmator controls
and address anything else that surfaces at that time.
WE hope that we can hand off the system to the next user (pcatryss::
Peter Catryssee late in the afternoon Tuesday and my reservation after
his on coral is to allow him time to write if we extend into his
I will be away from the lab from now until Tuesday but available by cell
Thank you for your support!
More information about the raith